OpenStudy (anonymous):

What is the standard deviation of the following data set? 11, 13, 15, 16, 20, 21 A) √15.2 B) √18 C) √76 D) √3.6

OpenStudy (anonymous):

c

OpenStudy (anonymous):

standard deviation is \[\sigma=\sqrt{1/n(\sum_{1}^{n}(x-x^{-})^2}\] where xbar is mean mean is=(11+13+15+16)20+21)6=16 ans standard deviation using above formula is =3.55

OpenStudy (anonymous):

d